The superconducting transition temperature in a metal is 4.2 K for zero magnetic
field. If the magnetic field required to destroy superconductivity at 3.2 K is 0.02T,
calculate the critical magnetic field at absolute zero.
1
Expert's answer
2019-03-05T09:41:19-0500

The critical magnetic field


"B_c(T)=B_c(0)\left[1-\left(\frac{T}{T_c}\right)^2\right]"

So,


"B_c(0)=\frac{B_c(T)}{\left[1-\left(\frac{T}{T_c}\right)^2\right]}=\frac{0.02}{\left[1-\left(\frac{3.2}{4.2}\right)^2\right]}=0.048 \;\rm{T}"